# Coarse-Grained Games: A Framework for Bounded Perception in Game Theory

### v1

@@ 7.2 Adverse Selection
In this analysis, we reinterpret [1]’s concept of “less information” through the lens of low resolution in a coarse-grained game. In the original discussion of adverse selection, the concept of the lemon market is examined. Adverse selection refers to a situation in which information asymmetry in a market causes lower-quality goods or higher-risk transactions to dominate, ultimately leading to a decline in market quality and efficiency.
@@
In this section, we reconstruct the lemon market within the framework of a coarse-grained game. We define the players and their strategies as follows:
@@
$\psi_{\text{EMP}}(\phi_{C}(u_{C}(s_{j_{C}})))-\text{sale price},\quad(s_{j_{C}}\text{ is ``Buy }m_{1}\text{'' or ``Buy }m_{2}\text{''})$ ” )
@@
sale $\text{sale price}-\psi_{\text{EMP}}(\phi_{D}(u_{D}(s_{j_{D}}))),\quad(s_{j_{D}}\text{ is ``Sell }m_{1}\text{'' or ``Sell }m_{2}\text{''}).$ ” ) .
@@
However, in a coarse-grained game, the amount $C$ is willing to pay depends on their coarse-grained partition. For instance, if $C$’s partition is such that they cannot distinguish between cars valued between \$10,000 and \$20,000, then, under EMP, $C$ evaluates both $m_{1}$ and $m_{2}$ as worth \$15,000. Consequently, $C$ signals their willingness to pay up to \$15,000 for either car.
@@
This conclusion aligns with the original insight from the lemon market model, where high-quality products exit the market, leaving only low-quality products in circulation. However, by adopting the framework of a coarse-grained game, we can provide a more realistic interpretation in the following ways:
@@
    A More Realistic Consumer Evaluation Process: In Akerlof’s original lemon market model, consumer $C$ makes purchasing decisions based on the market’s average price, which assumes an expectation over the distribution of product qualities. However, this assumption becomes unrealistic in extreme cases. For instance, when faced with both an exceptionally low-quality used car and a premium-quality used car, it is implausible for $C$ to simply rely on the market’s average price as a valuation. $C$ would likely recognize the flaws in an extremely low-quality used car and abandon probabilistic reasoning altogether. In a coarse-grained game, the consumer’s perception follows a predefined partition, which sets realistic limits on their ability to distinguish quality differences.
@@
    The Effect of High-Resolution Consumers ($C^{\prime}$) on the Market: If a high-resolution consumer $C^{\prime}$ enters the market—one who can correctly assess the value of a used car at \$20,000—the dealer $D$ would no longer be able to sell a lemon to them. Importantly, this phenomenon does not occur because $C^{\prime}$ has access to more external information, nor because $D$ has provided additional details. Instead, it happens because $C^{\prime}$ possesses a greater ability to interpret the same information accurately. This perspective is difficult to explain within the conventional framework of information asymmetry but can be naturally accounted for using the coarse-grained game approach.
@@ 8 Conclusion
In this paper, we introduced coarse-grained games as a framework for analyzing decision-making under perceptual constraints. While coarse-graining is a standard tool in the natural sciences, it has been largely absent from game-theoretic models in the social sciences. We argue that this omission is problematic because human decision-making is inherently coarse-grained; individuals simplify information when making economic, legal, and political choices.
@@
    In coarse-grained games, the distinction between subjective and objective perspectives, as well as their comparison, is meaningful. A subjective payoff matrix is obtained based on a player’s level of coarse perception—equivalently, their resolution—and can be viewed as a blurred version of the underlying objective payoff matrix.
@@
    In multi-player settings, where each player may have a different coarse-grained perception, the strategic outcomes can diverge from those derived from a game in which all players act based on the same objective payoff matrix. However, when players adopt entropy-maximizing preprocessing, the Nash equilibria of the coarse-grained game do not disappear or become entirely random. Instead, they are transformed in a systematic way. In particular, we highlight the following key results:
@@
      In non-repeated pure-strategy games, a Nash equilibrium in the original game remains a Nash equilibrium in the coarse-grained game. However, a strategy profile that was not a Nash equilibrium in the original game may become a Nash equilibrium in the coarse-grained game.
@@
      In non-repeated mixed-strategy games, if the level of coarse-graining is sufficiently high, the coarse-grained game degenerates into a simplified version of the original game.
@@
Furthermore, this paper demonstrates that coarse-grained games are applicable to the social sciences by examining two economic phenomena. The first concerns how consumers’ coarse-grained perceptions affect their behavior in response to minor model changes in products. A key finding is that when the benefit perceived by consumers from a minor model change is too small, they are unable to distinguish between the old and new versions of the product. This results in an equilibrium selection problem. In such cases, overly subtle model updates fail to generate quality-based competition and instead rely on external focal points—such as consumers’ preference for novelty or the persuasiveness of sales strategies—to influence purchasing decisions.
The second theme revisits the classic lemon market problem through the lens of coarse-grained games. The analysis reveals that under certain conditions, the market discourages the sale of high-quality used cars while promoting the sale of low-quality ones. However, this phenomenon is not necessarily due to an objective asymmetry of information, but rather to a subjective asymmetry in consumers’ ability to interpret information—that is, a disparity in cognitive resolution.
Together, these two examples illustrate not only that coarse-grained games can be applied to economic modeling, but also that their analytical outcomes often align closely with real-world intuition.
This paper has the following limitations. First, while coarse-grained games were defined using coarse set theory, we have not explored alternative formulations that incorporate more complex structural elements. Second, in mapping sets of real values back to single real numbers, we exclusively employed the entropy-maximizing strategy preprocessing (EMP). Further investigation is needed to assess how different strategy preprocessing methods influence the behavior of coarse-grained games. Finally, our discussion on applications to the social sciences was limited to the formulation of hypotheses rather than empirical validation. We leave these issues for future research.

### v2

@@ 7.2 Adverse Selection
In this analysis, we reinterpret [2]’s concept of “less information” through the lens of low resolution in a CGG. In the original discussion of adverse selection, the concept of the lemon market is examined. Adverse selection refers to a situation in which information asymmetry in a market causes lower-quality goods or higher-risk transactions to dominate, ultimately leading to a decline in market quality and efficiency.
@@
In this section, we reconstruct the lemon market within the framework of a CGG. We define the players and their strategies as follows:
@@
$\psi_{\text{EMP}}(\varphi_{C}(u_{C}(s_{j_{C}})))-\text{sale price},\quad(s_{j_{C}}\text{ is ``Buy }m_{1}\text{'' or ``Buy }m_{2}\text{''})$ ” )
@@
sale $\text{sale price}-\psi_{\text{EMP}}(\varphi_{D}(u_{D}(s_{j_{D}}))),\quad(s_{j_{D}}\text{ is ``Sell }m_{1}\text{'' or ``Sell }m_{2}\text{''}).$ ” ) .
@@
However, in a CGG, the amount $C$ is willing to pay depends on their coarse-grained partition. For instance, if $C$’s partition is such that they cannot distinguish between cars valued between \$10,000 and \$20,000, then, under EMP, $C$ evaluates both $m_{1}$ and $m_{2}$ as worth \$15,000. Consequently, $C$ signals their willingness to pay up to \$15,000 for either car.
@@
This conclusion aligns with the original insight from the lemon market model, where high-quality products exit the market, leaving only low-quality products in circulation. However, by adopting the framework of a CGG, we can provide a more realistic interpretation in the following ways:
@@
    A More Realistic Consumer Evaluation Process: In Akerlof’s original lemon market model, consumer $C$ makes purchasing decisions based on the market’s average price, which assumes an expectation over the distribution of product qualities. However, this assumption becomes unrealistic in extreme cases. For instance, when faced with both an exceptionally low-quality used car and a premium-quality used car, it is implausible for $C$ to simply rely on the market’s average price as a valuation. $C$ would likely recognize the flaws in an extremely low-quality used car and abandon probabilistic reasoning altogether. In a CGG, the consumer’s perception follows a predefined partition, which sets realistic limits on their ability to distinguish quality differences.
@@
    The Effect of High-Resolution Consumers ($C^{\prime}$) on the Market: If a high-resolution consumer $C^{\prime}$ enters the market—one who can correctly assess the value of a used car at \$20,000—the dealer $D$ would no longer be able to sell a lemon to them. Importantly, this phenomenon does not occur because $C^{\prime}$ has access to more external information, nor because $D$ has provided additional details. Instead, it happens because $C^{\prime}$ possesses a greater ability to interpret the same information accurately. This perspective is difficult to explain within the conventional framework of information asymmetry but can be naturally accounted for using the CGG approach.
@@ 8 Conclusion
In this paper, we introduced Coarse-Grained Games (CGGs) as a framework for analyzing decision-making under perceptual constraints. While coarse-graining is a standard tool in the natural sciences, it has been largely absent from game-theoretic models in the social sciences. We argue that this omission is problematic because human decision-making is inherently coarse-grained; individuals simplify information when making economic, legal, and political choices.
@@
    In CGGs, the distinction between subjective and objective perspectives, as well as their comparison, is meaningful. A subjective payoff matrix is obtained based on a player’s level of coarse perception—equivalently, their resolution—and can be viewed as a blurred version of the underlying objective payoff matrix.
@@
    In multi-player settings, where each player may have a different coarse-grained perception, the strategic outcomes can diverge from those derived from a game in which all players act based on the same objective payoff matrix. However, when players adopt entropy-maximizing preprocessing, the Nash equilibria of the CGG do not disappear or become entirely random. Instead, they are transformed in a systematic way. In particular, we highlight the following key results:
@@
      In non-repeated pure-strategy games, a Nash equilibrium in the original game remains a Nash equilibrium in the CGG. However, a strategy profile that was not a Nash equilibrium in the original game may become a Nash equilibrium in the CGG.
@@
      In non-repeated mixed-strategy games, if the level of coarse-graining is sufficiently high, the CGG degenerates into a simplified version of the original game.
@@
Furthermore, this paper demonstrates that CGGs are applicable to the social sciences by examining two economic phenomena. The first concerns how consumers’ coarse-grained perceptions affect their behavior in response to minor model changes in products. A key finding is that when the benefit perceived by consumers from a minor model change is too small, they are unable to distinguish between the old and new versions of the product. This results in an equilibrium selection problem. In such cases, overly subtle model updates fail to generate quality-based competition and instead rely on external focal points—such as consumers’ preference for novelty or the persuasiveness of sales strategies—to influence purchasing decisions.
The second theme revisits the classic lemon market problem through the lens of CGGs. The analysis reveals that under certain conditions, the market discourages the sale of high-quality used cars while promoting the sale of low-quality ones. However, this phenomenon is not necessarily due to an objective asymmetry of information, but rather to a subjective asymmetry in consumers’ ability to interpret information—that is, a disparity in cognitive resolution.
Together, these two examples illustrate not only that CGGs can be applied to economic modeling, but also that their analytical outcomes often align closely with real-world intuition.
This paper has the following limitations. First, while CGGs were defined using coarse set theory, we have not explored alternative formulations that incorporate more complex structural elements. Second, in mapping sets of real values back to single real numbers, we exclusively employed the entropy-maximizing strategy preprocessing (EMP). Further investigation is needed to assess how different strategy preprocessing methods influence the behavior of CGGs. Finally, our discussion on applications to the social sciences was limited to the formulation of hypotheses rather than empirical validation. We leave these issues for future research.